OpenStudy (anonymous):

1 5 % × 10 = _____ Select one of the options below as your answer: A. .02 B. .05 C. .15 D. 2

OpenStudy (oaktree):

Is that 15% of 10?

OpenStudy (goformit100):

D. 2

OpenStudy (anonymous):

remember that % means out of 100 so 5%= 5/100 15%=15/100

OpenStudy (oaktree):

Actually, that's wrong, @goformit100

OpenStudy (anonymous):

1/5=0.2 0.2/100=0.02 0.02*10=0.2

OpenStudy (anonymous):

@goformit100 20% of 10 is 2

OpenStudy (anonymous):

@DamnImBad 5%= .05 15%=.15 50%=.50 see the pattern